# Appearance-based Refinement for Object-Centric Motion Segmentation

### v1

@@ Abstract
The goal of this paper is to discover, segment, and track independently moving objects in complex visual scenes. Previous approaches have explored the use of optical flow for motion segmentation, leading to imperfect predictions due to partial motion, background distraction, and object articulations and interactions. To address this issue, we introduce an appearance-based refinement method that leverages temporal consistency in video streams to correct inaccurate flow-based proposals. Our approach involves a simple selection mechanism that identifies accurate flow-predicted masks as exemplars, and an object-centric architecture that refines problematic masks based on exemplar information. The model is pre-trained on synthetic data and then adapted to real-world videos in a self-supervised manner, eliminating the need for human annotations. Its performance is evaluated on multiple video segmentation benchmarks, including DAVIS, YouTubeVOS, SegTrackv2, and FBMS-59. We achieve competitive performance on single-object segmentation, while significantly outperforming existing models on the more challenging problem of multi-object segmentation. Finally, we investigate the benefits of using our model as a prompt for a per-frame Segment Anything Model.
Figure 1: Overview of our multi-object motion segmentation method. The method starts with proposing object masks based on optical flow inputs (i.e., flow-based proposals), followed by an appearance-based refinement of flow-predicted masks. The latter stage relies on a selection-correction procedure, where high-quality exemplar masks are selected to guide the refinement of other masks.
1. Introduction
@@
To address the limitations of flow-only object discovery, we draw inspiration from the two-stream hypothesis in neuroscience [70, 24, 60], which suggests that the human visual cortex processes high temporal frequency motion information in the dorsal stream, and slower temporal frequency appearance information in the ventral stream. This two-stream (motion & appearance) approach has been successfully applied to tasks such as action recognition [74, 22, 53, 23] and video object segmentation [20, 98, 103, 97, 100, 59, 48]. In this paper, we propose a two-stream approach for the task of object discovery, segmentation, and tracking in videos. The motion stream is used for object discovery, while the appearance stream is used for consistency in segmentation and tracking, leveraging long-term correlations that arise from the persistence of appearance.
@@
Our proposed approach is depicted in Figure 1 and consists of two stages. First, we use an off-the-shelf motion segmentation model to discover objects that move independently of the camera in some frames of the video sequence. This model generates proposals for segmentation masks in each frame, which vary in quality due to indistinct and partially occluded object motion. In the second stage, we leverage the temporally coherent appearance in videos for mask selection and mask correction. Mask selection involves picking high-quality proposal masks based on both temporal coherence across mask shapes and semantic consistency with object appearances. Mask correction involves using the selected masks (termed “exemplars”) to guide an appearance-based mask correction process in an object-centric model. Both selection and correction employ the self-supervised DINO [9] visual features. This selection-correction procedure enables accurate mask predictions under complex motion scenarios, including partial object motion, background distraction, object articulation, and interaction, etc.
The entire model is first pre-trained on synthetic videos, which enables it to be directly applied to real video sequences. To further bridge the domain gap between synthetic and real videos, we propose a self-supervised training scheme that serves as a simple yet effective Sim2Real adaptation procedure, further enhancing the model performance on real-world data. We evaluate the model on multiple public video object segmentation benchmarks, including DAVIS, YouTubeVOS, SegTrackv2, and FBMS-59. The model performs competitively in discovering and tracking single objects, while significantly outperforming existing approaches under more challenging scenarios involving multiple objects. Finally, we use our video level predictions as prompts for a Segment Anything Model (SAM) [38], and demonstrate that the object-centric consistency surpasses direct per-frame segmentation, also highlighting the complementary improvements that SAM could bring to our results.
2. Related work
@@
Recent research [20, 3, 99] has also focused on incorporating appearance information with motion cues. Choudhury et al. [14] and Ponimatkin et al. [66] leverage spectral clustering on appearance features to improve segmentation accuracy. DyStaB [97] trains dynamic and static models via an iterative bootstrapping strategy. Our method follows a similar idea and targets a more challenging multi-object motion scenario.
Video object segmentation focuses on obtaining pixel-level masks of one or multiple salient objects in videos. As defined by several benchmark datasets [65, 8], video object segmentation (VOS) tasks can be further divided into two main protocols, namely semi-supervised VOS [85, 40, 41, 59, 87, 30, 9] and unsupervised VOS [51, 84, 11, 45, 52]. The former provides first-frame ground-truth annotation and aims to recover segmentation masks for the remaining frames, while the latter automatically discovers and tracks salient objects from raw videos. Albeit referred to as “unsupervised”, some previous studies rely on models that are pre-trained using manual labels, whereas our proposed method does not rely on any human annotations.
@@
Early unsupervised object-centric learning research mainly focuses on synthetic datasets, while more recent works achieve object segmentation and tracking on real-world videos. Fan et al. [21] and Aydemir et al. [2] achieve an object-centric video representation by employing reconstruction of the self-supervised visual features, thereby extending the method of Seitzer et al. [71] from real-world images to videos. In this paper, we combine the feature reconstruction idea with motion priors to obtain object-centric representations in real-world videos.
Figure 2: Appearance-based mask selection. Top: The mask selector takes in RGB frames, DINO features, and flow-predicted masks, generating three-channel error maps for each object (in this case the standing person). The error maps are used to select the exemplar masks for each object. The false positive (FP) channel highlights the over-segmented areas (the dog on the right, as that is not part of the man), while the false negative (FN) channel predicts the missing masks (the lower body of the person). The remaining channel encompasses a combination of true positive and true negative (TP + TN) regions, where the flow-predicted masks align accurately with the groundtruth, as indicated by white regions. The groundtruth annotation boundaries (purple contours) are provided for illustration purposes only. Bottom: Two potential cues underlying the mask selection process: (i) temporal consistency of mask shapes across consecutive frames, and and (ii) the consistency of semantic content – the mask (including all its parts) should only correspond to one semantic class. In this case, the mask in the right frame is not selected, as it consists of two parts that separately correspond to different semantic classes (the person and the dog).
Figure 3: Appearance-based mask corrector. The object-centric architecture first utilises the exemplar mask information to initialise independent object queries. The transformer decoder module then refines the query vectors with two cross-attention layers using dense DINO features for one, and exemplar masks for the other. There are two major outputs: (i) the refined queries are spatially broadcast and projected to recover DINO features in target frames, and (ii) the cross-attention maps between queries and target frame DINO features are extracted, followed by a CNN-based upsampling module to yield final refined masks. Note, “Exemplar mask frames” refer to frames from which the exemplar masks are selected. For illustration, we present exemplar masks for different objects in the same frame. However, in practice, different objects may have distinct sets of exemplar mask frames.
3. Appearance-based selection and correction
@@ 3.1. Problem scenario
{^$\{\hat{\mathcal{M}}_{1}^{k},\dots,\hat{\mathcal{M}}_{T}^{k}\}_{k=1}^{K}=\Phi_{\text{OCLR-flow}}(\mathcal{V}_{\text{flow}})\vspace{-0.05cm}$ (1)
@@
Our goal is to refine video object segmentation by leveraging the video appearance stream. We follow a two-stage procedure that involves selecting a set of high-quality exemplar segmentation masks and refining them by exploiting the temporal consistency in the appearance stream. To achieve this, we use the flow-predicted masks as a given prior and apply the following model:
{~$,\displaystyle\{\tilde{\mathcal{M}}_{1}^{k},$~$\dots,\tilde{\mathcal{M}}_{T}^{k}\}_{k=1}^{K}=$
$\Phi_{\text{corrector}}\circ\Phi_{\text{selector}}(\mathcal{V}_{\text{RGB}};\{\hat{\mathcal{M}}_{1}^{k},\dots,\hat{\mathcal{M}}_{T}^{k}\}_{k=1}^{K})$ (2)
@@
3.2. Appearance-based mask selection
@@
We propose a mask selector that takes the flow-predicted masks and RGB frames as inputs, to identify a set of high-quality exemplar masks for each object $k$ independently. Formally,
(^$(\hat{\mathcal{E}}_{1}^{k},\dots,\hat{\mathcal{E}}_{T}^{k})=\Phi_{\text{selector}}($ $\mathcal{V}_{\text{RGB}};\hat{\mathcal{M}}_{1}^{k},\dots,\hat{\mathcal{M}}_{T}^{k})$
[…]
@@
where ^$\hat{\mathcal{E}}_{1}^{k},\dots,\hat{\mathcal{E}}_{T}^{k}$ denote the inferred error maps as intermediate results, from which a set of $t$ exemplar masks $\hat{\mathcal{M}_{e}}_{1}^{k},\dots,\hat{\mathcal{M}_{e}}_{t}^{k}$ can be determined for each object by comparing the size of predicted erroneous regions.
@@
By considering multiple frames and their associated flow-based mask proposals, a selection procedure is able to employ two potential cues, as illustrated in Figure 2: (i) the temporal consistency between neighbouring masks – objects undergo smooth and continuous changes in position and shape, and the selector can filter out the frames with empty or incomplete masks due to stationary objects (i.e., no flow) or abrupt occlusions; (ii) the consistency of semantic content within corresponding RGB frames – as high-quality object mask normally corresponds to only one semantic class and this does not switch through the video.
@@
Architecture. Figure 2 illustrates a simple architecture for the mask selector that considers a joint input of RGB frames and flow-predicted masks at all time steps. To incorporate semantic priors, visual features from the frozen DINO ViT [9] are concatenated channel-wise with the CNN-encoded features in the model bottleneck. These DINO features have been shown to be effective for encoding frame-to-frame correspondence and semantic groupings [73, 56, 89, 72], and thus enable the transformer to use the two potential cues for the selection. Full-resolution error maps are obtained from a UNet-like upsampling process, where the original RGB frames help by providing sharp boundaries.
@@
Training objective. We task the selector to predict an error map, i.e., the difference map between the input flow-predicted mask and the groundtruth segmentation. Specifically, for each frame $i$, the prediction ^$\hat{\mathcal{E}}_{i}\in\mathbb{R}^{H\times W\times 3}$ consists of three channels corresponding to the false positive (FP), false negative (FN) and a combination of true positive and true negative (TP + TN) regions. We train the mask selector by applying a binary cross-entropy loss on each of the predicted regions:
$\mathcal{L}_{\text{selector}}=\frac{1}{T}\sum_{i=1}^{T}\left(\mathcal{L}_{\text{bce}}(\hat{\mathcal{E}}_{i},\mathcal{E}_{i})\right)\vspace{-0.05cm}$ (4)
@@
The groundtruth error map, $\mathcal{E}_{i}$, is derived from a synthetic video dataset (as will be detailed in Sec. 3.4. Sim2Real training procedure), which serves as the training data for the model. By comparing the size of erroneous regions (i.e., FP and FN regions) across the sequence, we can compute a score that indicates the mask quality for each frame, and subsequently, we keep a set of $t$ exemplar masks with high scores for each object.
3.3. Appearance-based mask correction
@@
{~$\{\tilde{\mathcal{M}}_{1}^{k},\dots,\tilde{\mathcal{M}}_{T}^{k}\}_{k=1}^{K}=$
$\Phi_{\text{corrector}}(\mathcal{V}_{\text{RGB}};\{\hat{\mathcal{M}}_{1}^{k},\dots,\hat{\mathcal{M}}_{T}^{k};{\text{\small$\hat{\mathcal{M}_{e}}$}_{1}^{k},\dots,\text{\small$\hat{\mathcal{M}_{e}}$}_{t}^{k}}\}_{k=1}^{K})$ (5)
@@
In the first cross-attention module, the query vectors attend to the visual features of all target frames, whereas in the second cross-attention step, the key and value entries are generated by combining exemplar masks with the visual features of corresponding frames.
@@
$L_{\text{mask}}=\frac{1}{KT}\sum_{i=1}^{T}\sum_{k=1}^{K}\Big{(}\Big{.}$ $\lambda_{\text{bce}}L_{\text{bce}}(\tilde{\mathcal{M}}_{i}^{k},{\mathcal{M}}_{i}^{k})$
$+$ $\lambda_{\text{bound}}L_{\text{bound}}(\tilde{\mathcal{M}}_{i}^{k},{\mathcal{M}}_{i}^{k})\Big{.}\Big{)}\vspace{-0.1cm}$ (6)
with $\lambda_{\text{bce}},\lambda_{\text{bound}}$ denoting weight factors. The groundtruth segmentation masks $\{{\mathcal{M}}_{i}^{k}\}$ are provided by the synthetic video sequences, as will be detailed in Sec. 3.4. Sim2Real training procedure.
@@
$L_{\text{feat-recon}}=\frac{1}{T}\sum_{i=1}^{T}\Big{(}\Big{.}$ $\lambda_{\text{recon}}\text{MSE}(\hat{d}_{i},d_{i})$
$+$ $\frac{1}{K}\sum_{k=1}^{K}\lambda_{\text{alpha}}L_{\text{bce}}(\hat{\alpha}_{i}^{k},\alpha_{i}^{k})\Big{.}\Big{)}$ (7)
@@
3.4. Sim2Real training procedure

### v2

@@ Abstract
The goal of this paper is to discover, segment, and track independently moving objects in complex visual scenes. Previous approaches have explored the use of optical flow for motion segmentation, leading to imperfect predictions due to partial motion, background distraction, and object articulations and interactions. To address this issue, we introduce an appearance-based refinement method that leverages temporal consistency in video streams to correct inaccurate flow-based proposals. Our approach involves a sequence-level selection mechanism that identifies accurate flow-predicted masks as exemplars, and an object-centric architecture that refines problematic masks based on exemplar information. The model is pre-trained on synthetic data and then adapted to real-world videos in a self-supervised manner, eliminating the need for human annotations. Its performance is evaluated on multiple video segmentation benchmarks, including DAVIS, YouTubeVOS, SegTrackv2, and FBMS-59. We achieve competitive performance on single-object segmentation, while significantly outperforming existing models on the more challenging problem of multi-object segmentation. Finally, we investigate the benefits of using our model as a prompt for the per-frame Segment Anything Model.
Keywords:
Motion Segmentation Video Object Segmentation
1 Introduction
Figure 1: Overview of our multi-object motion segmentation method. The method starts with proposing object masks based on optical flow inputs (i.e., flow-based proposals), followed by an appearance-based refinement of flow-predicted masks. The latter stage relies on a selection-correction procedure, where high-quality exemplar masks are selected to guide the correction of other masks. Mask selection involves picking high-quality proposal masks based on both temporal coherence across mask shapes and semantic consistency with object appearances. Mask correction involves using the selected masks (the exemplars) to guide an appearance-based mask correction process in an object-centric model. The success of the method can be seen in the bottom row where correct segmentation masks are predicted, overcoming the deficiencies of the flow-predicted masks in Stage 1.
@@ 1 Introduction
In this paper, to address the drawbacks of flow, we incorporate temporally-consistent appearance information by using a two-stream approach [82, 26, 23] for the task of object discovery, segmentation, and tracking in videos. The motion stream (flow) is used for object discovery, while the appearance stream is used for maintaining consistency in segmentation and tracking, leveraging long-term correlations that arise from the persistence of appearance.
@@
Our approach is depicted in Figure 1 and consists of two stages. In the first stage, we use an off-the-shelf flow-based motion segmentation model, OCLR [102]. As noted above, the flow can be used to discover moving objects (and has little Sim2Real gap), but this is as far as OCLR-flow can go – the model generates proposals for segmentation masks in each frame, but these vary in quality due to indistinct and partially occluded object motion. Appearance can give stronger segmentations, yet faces a significant Sim2Real gap. The novelty in this paper lies in our solution to this problem – “How can appearance information be used to improve flow-based segmentation?” – with three distinct innovations that form the second stage of the approach: (i) we propose a sequence-level mask selection mechanism, with DINO features [11] employed for establishing temporal consistency; (ii) the selected masks (that we treat as exemplars) support mask correction, where we employ a combination of direct mask supervision and motion-initialised semantic reconstruction for each frame; (iii) a self-supervised adaptation technique is introduced to effectively address the Sim2Real gap (that has some similarities with [94]). This selection-correction procedure enables accurate mask predictions under complex motion scenarios, including partial object motion, background distraction, object articulation, and interaction, etc.
In summary, we make the following five contributions:
  1. 1.
    We introduce a sequence-level selection mechanism designed to globally identify exemplar object masks from flow-based segmentation predictions over a long temporal range.
  2. 2.
    We propose an object-centric mask correction network that incorporates both local (i.e., short-term) and global (i.e., exemplar) information into motion-initialised object queries. Additionally, we investigate a unique loss formulation that combines direct mask supervision with motion-aware semantic reconstruction.
  3. 3.
    We follow a Sim2Real training procedure that relies on zero human annotation: The framework is synthetically pre-trained, followed by a self-supervised training scheme designed for simple yet effective domain adaptation to real-world video sequences.
  4. 4.
    Our model is evaluated on multiple motion segmentation benchmarks, including DAVIS, YouTubeVOS, SegTrackv2, and FBMS-59, demonstrating competitive performance on single-object segmentation, while significantly outperforming existing approaches under challenging multi-object scenarios.
  5. 5.
    In further exploration, we use our video-level predictions as prompts for the Segment Anything Model (SAM) [43]. This reveals that the object-centric consistency surpasses direct per-frame segmentation, also highlighting the complementary improvements that SAM could bring to our results.
2 Related Work
@@
Recent research [23, 3, 109] has also focused on incorporating appearance information with motion cues. Choudhury et al. [16] and Ponimatkin et al. [73] leverage spectral clustering on appearance features to improve segmentation accuracy. DyStaB [106] trains dynamic and static models via an iterative bootstrapping strategy. Our method follows a similar two-stream idea and targets a more challenging multi-object motion scenario.
Video Object Segmentation (VOS) focuses on obtaining pixel-level masks of one or multiple salient objects in videos. As defined by several benchmark datasets [71, 10], video object segmentation tasks can be further divided into two main protocols, namely semi-supervised VOS [93, 47, 46, 65, 96, 34, 11] and unsupervised VOS [57, 92, 13, 51, 58]. The former provides first-frame ground-truth annotation and aims to recover segmentation masks for the remaining frames, while the latter automatically discovers and tracks salient objects from raw videos. Albeit referred to as “unsupervised”, some previous studies rely on models that are pre-trained using manual labels, whereas our proposed method does not rely on any human annotations.
Sequence-Level Frame Selection aims to pick the most informative frames from a video sequence, enhancing efficiency in downstream applications. This idea is widely explored across various computer vision domains, including action recognition [76, 116, 75], video retrieval [114, 8], etc. In video segmentation, BubbleNet [30] employs this mechanism to select a single RGB frame that initiates mask propagation for semi-supervised VOS, while Yin et al. [110] extends this idea to interactive VOS. Wu et al. [100] introduce a Memory Selection Network for memory-based mask propagation. Recent studies [45, 18, 5] also utilise temporal selection to propose key frames for efficient segmentation annotations. Unlike prior works that focus on RGB frame selections via confidence score estimations, our selection network regards error map predictions as a proxy task to guide the exemplar mask selection for each object.
Two-Stream Approaches. The idea originates from the two-stream hypothesis in neuroscience [78, 27, 66], which suggests that the human visual cortex processes high temporal frequency motion information in the dorsal stream, and slower temporal frequency appearance information in the ventral stream. This two-stream (motion & appearance) approach has been successfully applied to tasks such as action recognition [82, 26, 59, 25] and VOS [23, 108, 115, 106, 111, 65, 54].
@@
Early unsupervised object-centric learning research mainly focuses on synthetic datasets, while more recent works achieve object segmentation and tracking on real-world videos. Fan et al. [24] and Aydemir et al. [2] achieve an object-centric video representation by employing reconstruction of the self-supervised visual features, thereby extending the method of Seitzer et al. [79] from real-world images to videos.
3 Appearance-Based Selection and Correction
@@ 3.1 Problem Scenario
{^$\{\hat{\mathcal{M}}_{1}^{k},\dots,\hat{\mathcal{M}}_{T}^{k}\}_{k=1}^{K}=\Phi_{\text{OCLR-flow}}(\mathcal{V}_{\text{flow}})$ (1)
@@
Our goal is to refine motion segmentation by leveraging the video appearance stream, specifically, it involves selecting a set of high-quality exemplar segmentation masks and refining them by exploiting the temporal consistency in the appearance stream. To achieve this, we use the flow-predicted masks as a given prior and apply the following model:
{~$\{\tilde{\mathcal{M}}_{1}^{k},\dots,\tilde{\mathcal{M}}_{T}^{k}\}_{k=1}^{K}=\Phi_{\text{corrector}}\circ\Phi_{\text{selector}}(\mathcal{V}_{\text{RGB}};\{\hat{\mathcal{M}}_{1}^{k},\dots,\hat{\mathcal{M}}_{T}^{k}\}_{k=1}^{K})$ (2)
@@
3.2 Appearance-Based Mask Selection
Figure 2: Appearance-based mask selection. Left: The mask selector takes in RGB frames, DINO features, and flow-predicted masks, generating three-channel error maps for each object (in this case the standing person). The error maps are used to select the exemplar masks for each object. The false positive (FP) channel highlights the over-segmented areas (the dog), while the false negative (FN) channel predicts the missing masks (the lower body of the person). The remaining channel encompasses a combination of true positive and true negative (TP + TN) regions, where the flow-predicted masks align accurately with the groundtruth, as indicated by white regions. The groundtruth annotation boundaries (purple contours) are provided for illustration purposes only. Right: Two potential cues underlying the mask selection process: (i) temporal consistency of mask shapes across consecutive frames, and and (ii) the consistency of semantic content – the mask (including all its parts) should only correspond to one semantic class.
@@ 3.2 Appearance-Based Mask Selection
We propose a mask selector that takes the flow-predicted masks and RGB frames as inputs, to identify a set of high-quality exemplar masks independently for each object $k$. Formally,
(^$(\hat{\mathcal{E}}_{1}^{k},\dots,\hat{\mathcal{E}}_{T}^{k})=\Phi_{\text{selector}}(\mathcal{V}_{\text{RGB}};\hat{\mathcal{M}}_{1}^{k},\dots,\hat{\mathcal{M}}_{T}^{k}){\;\longrightarrow(\text{\small$\hat{\mathcal{M}_{e}}$}_{1}^{k},\dots,\text{\small$\hat{\mathcal{M}_{e}}$}_{t}^{k})}$ (3)
@@
where ^$\hat{\mathcal{E}}_{1}^{k},\dots,\hat{\mathcal{E}}_{T}^{k}$ are predicted error maps, from which a set of $t$ exemplar masks $\hat{\mathcal{M}_{e}}_{1}^{k},\dots,\hat{\mathcal{M}_{e}}_{t}^{k}$ can be determined for each object by comparing the size of predicted erroneous regions.
@@
By considering multiple frames and their associated flow-based mask proposals, the selection procedure is able to employ two potential cues, as illustrated in Figure 2: (i) the temporal consistency between neighbouring masks – objects undergo smooth and continuous changes in position and shape, and the selector can filter out the frames with empty or incomplete masks due to stationary objects (i.e., no flow) or abrupt occlusions; (ii) the consistency of semantic content within each RGB frame – as a high-quality object mask normally corresponds to only one semantic class and this does not switch throughout the video.
@@
Architecture. Figure 2 illustrates a simple architecture for the mask selector that considers a joint input of RGB frames and flow-predicted masks at all time steps. To incorporate semantic priors, visual features from the frozen DINO ViT [11] are concatenated channel-wise with the CNN-encoded features in the model bottleneck. These DINO features have been shown to be effective for encoding frame-to-frame correspondence [81, 62, 98, 80], and thus enable the transformer to use the two potential cues for the selection. Full-resolution error maps are obtained from a UNet-like upsampling process, where the original RGB frames help by providing sharp boundaries.
@@
Training Objective. In contrast to prior approaches [30, 110] that directly compute confidence scores, we task the selector to predict an error map, i.e., the difference map between the input flow-predicted mask and the groundtruth segmentation, as an intermediate result. Specifically, for each frame $i$, the prediction ^$\hat{\mathcal{E}}_{i}\in\mathbb{R}^{H\times W\times 3}$ consists of three channels corresponding to the false positive (FP), false negative (FN) and a combination of true positive and true negative (TP + TN) regions. We train the mask selector by applying a binary cross-entropy loss on each of the predicted regions:
$\mathcal{L}_{\text{selector}}=\frac{1}{T}\sum_{i=1}^{T}\left(\mathcal{L}_{\text{bce}}(\hat{\mathcal{E}}_{i},\mathcal{E}_{i})\right)$ (4)
@@
The groundtruth error map, $\mathcal{E}_{i}$, is derived from a synthetic video dataset (as will be detailed in Sec. 3.4), which serves as the training data for the model. By comparing the size of erroneous regions (i.e., FP and FN regions) across the sequence, we can compute a score that indicates the mask quality for each frame, and subsequently keep $t$ exemplar masks with high scores for each object.
3.3 Appearance-Based Mask Correction
Figure 3: Appearance-based mask corrector. The object-centric architecture first utilises the exemplar mask information to initialise independent object queries. The transformer decoder module then refines the query vectors with two cross-attention layers using dense target frame features for one, and exemplar information for the other. There are two major outputs: (i) the refined queries are spatially broadcast and projected to recover DINO features in target frames, and (ii) the cross-attention maps between queries and target frame DINO features are extracted, followed by a CNN-based upsampling module to yield final refined masks. Note, “Exemplar mask frames” refer to frames from which the exemplar masks are selected. For illustration, we present exemplar masks for different objects in the same frame. However, in practice, different objects may have distinct sets of exemplar mask frames.
@@ 3.3 Appearance-Based Mask Correction
{~$\{\tilde{\mathcal{M}}_{1}^{k},\dots,\tilde{\mathcal{M}}_{T}^{k}\}_{k=1}^{K}=\Phi_{\text{corrector}}(\mathcal{V}_{\text{RGB}};\{\hat{\mathcal{M}}_{1}^{k},\dots,\hat{\mathcal{M}}_{T}^{k};{\text{\small$\hat{\mathcal{M}_{e}}$}_{1}^{k},\dots,\text{\small$\hat{\mathcal{M}_{e}}$}_{t}^{k}}\}_{k=1}^{K})$ (5)
@@
In the first cross-attention layer, the query vectors attend to the visual features of all adjacent target frames, whereas in the second cross-attention layer, the key and value entries are generated by combining exemplar masks with the visual features of corresponding frames. These layers enable the incorporation of short-term local information and global exemplar information, respectively.
@@
$L_{\text{mask}}=\frac{1}{KT}\sum_{i=1}^{T}\sum_{k=1}^{K}\Big{(}\Big{.}\lambda_{\text{bce}}L_{\text{bce}}(\tilde{\mathcal{M}}_{i}^{k},{\mathcal{M}}_{i}^{k})+\lambda_{\text{bound}}L_{\text{bound}}(\tilde{\mathcal{M}}_{i}^{k},{\mathcal{M}}_{i}^{k})\Big{.}\Big{)}$ (6)
with $\lambda_{\text{bce}},\lambda_{\text{bound}}$ denoting weight factors. The groundtruth segmentation masks $\{{\mathcal{M}}_{i}^{k}\}$ are provided by synthetic video sequences, as will be detailed in Sec. 3.4.
@@
$L_{\text{feat-recon}}=\frac{1}{T}\sum_{i=1}^{T}\Big{(}\Big{.}\lambda_{\text{recon}}\text{MSE}(\hat{d}_{i},d_{i})+\frac{1}{K}\sum_{k=1}^{K}\lambda_{\text{alpha}}L_{\text{bce}}(\hat{\alpha}_{i}^{k},\alpha_{i}^{k})\Big{.}\Big{)}$ (7)
@@
3.4 Sim2Real Training Procedure